Grade 11Magical Mathematics[Interesting Approach]

The value of k for which the expression x2+2xy+ky2 +2x+k = 0 can be resolved into 2 linear factors is given by?

Ranjita yadav
14 Years agoGrade 11

6 Answers

nikhil arora
ApprovedApproved Tutor Answer14 Years ago

this is an equation of second degree. if this equation can be resolved into linear factors, the linear factors will represent two straight lines which means the given equation should be pair of straight lines.

for pair of straight lines, Δ =0

comparing given equation with ax2 + 2hxy + by2 + 2gx + 2fy + c =0

putting  Δ = abc+2fgh-af2-bg2-ch2 = 0

k2-2k = 0

k(k-2)=0

k=0 or k=2

 

regards

Nikhil

Ranjita yadav
14 Years ago

can u post the ans. considering 11th syllabus.. coz I dnt thnk its der in 11th..!!

nikhil arora
14 Years ago

i posted answer with respect to context of coordinate geometry, and its there in 11th syllabus. if u are confused on Δ then its very useful concept , u must learn it.

Ranjita yadav
14 Years ago

can u explain that concept or give any link where I can get to know about this...??

nikhil arora
14 Years ago

general equation of second degree is given as

ax2 + 2hxy + by2 + 2gx + 2fy + c =0

this equation may represent any conic section under different conditions. so, we define Δ here.

Δ = abc+2fgh-af2-bg2-ch2 = 0

 

Conditions equation of the conic represents

Δ=0                                                    pair of straight lines

h=0 , a=b                                              circle

Δ≠0 ,h2-ab=0                                          parabola

Δ≠0, h2-ab<0                                             ellipse 

Δ≠0, h2-ab>0                                             hyperbola

Δ≠0, h2-ab>0 , a+b=0                                   rectangular hyperbola

Ranjita yadav
14 Years ago

Thanx a lot for ur help..!!